a coil of and radius 2 cm carrying a current at a rate of 1.9A/s is placed coaxially inside a solenoid of 500 turns and having radius 5 cm. find the induced emf produced in the solenoida coil of and radius 2 cm carrying a current at a rate of 1.9A/s is placed coaxially inside a solenoid of 500 turns and having radius 5 cm. find the induced emf produced in the solenoid
1
Expert's answer
2020-06-25T09:36:07-0400

Assume that "N_2=200"


"B_1=\mu_0\frac{N_1i_1}{2r_1}"


Magnetic flux linked with coil is given by


"\phi_2=N_2B_1A_2=N_2\mu_0\frac{N_1i_1}{2r_1}A_2"


According to definition of mutual inductance


"\phi_2=Mi_1 \to N_2\mu_0\frac{N_1i_1}{2r_1}A_2=Mi_1 \to M=N_2\mu_0\frac{N_1}{2r_1}A_2="


"=N_2\mu_0\frac{N_1}{2r_1}\cdot\pi r_2^2=200\cdot4\pi \cdot10^{-7}\cdot\frac{500}{2\cdot0.05}\cdot 3.14\cdot 0.02^2=0.00158H"


"emf=M\frac{di}{dt}=0.00158\cdot 1.9\approx3mV"